OpenStudy (anonymous):

Toricelli's theorem states that if there is a hole in a container of liquid h feet below the surface of the liquid, the the liquid will flow out at a rate given by R(h)= sort(ugh) where g = 32 ft/sec. Find a linear function that can be used to approximate this rate for holes that are close to 25 ft below the surface of the water

OpenStudy (anonymous):

*Torricelli's *then *sqrt(rgh)

OpenStudy (anonymous):

@SolomonZelman @FibonacciChick666 @DanJS

OpenStudy (anonymous):

sorry for all the typos btw :/

OpenStudy (fibonaccichick666):

lol I thought they were funny

OpenStudy (danjs):

just use the valuses they gave you

OpenStudy (anonymous):

its real that simple?

OpenStudy (fibonaccichick666):

you might like to read this: http://tutorial.math.lamar.edu/Classes/CalcI/LinearApproximations.aspx

OpenStudy (surry99):

and Torricelli's theorem says ...velocity = (2*g*h)^1/2

OpenStudy (solomonzelman):

I believe it is, \(\large\color{black}{ L(h) \approx R(25)+ R'(25)(h-25) }\), but I can be wrong.

OpenStudy (fibonaccichick666):

but yes, all of the above first, plug in your values given then apple

OpenStudy (anonymous):

okay! i'll read the notes… hang on. We did linear approximation during the first weeks of the smemester so I'm a bit foggy on it now; Im studying for my final

OpenStudy (fibonaccichick666):

yea, but do we know r?

OpenStudy (anonymous):

i don't believe so; they just gave me g

OpenStudy (solomonzelman):

it says 25 ft below the surface?

OpenStudy (fibonaccichick666):

you are given this ya? \[R(h)=\sqrt{rgh}\]

OpenStudy (fibonaccichick666):

we have to define r first, before we can just plug in

OpenStudy (fibonaccichick666):

Which as solomon pointed out, is done by reading between the lines in the wuetion

OpenStudy (anonymous):

oh shoot! no ! its R(h)= square root(2gh)

OpenStudy (fibonaccichick666):

that makes a difference

OpenStudy (anonymous):

it sure does

OpenStudy (fibonaccichick666):

and makes it easier haha

OpenStudy (fibonaccichick666):

ok so can you find \[R'(h)=?\]

OpenStudy (fibonaccichick666):

remember we have \[\sqrt{2*32*h}=R(h)\]

OpenStudy (anonymous):

so i just find the derivative and then do the whole linear approximation equation..? or am i missing something

OpenStudy (fibonaccichick666):

nope, that's it

OpenStudy (anonymous):

oh okay! thats not bad at all :) i always freak out when i see word problems and my mind freezes and forgets everything i know haha lol

OpenStudy (fibonaccichick666):

just remember to relax, write your givens, what you want, then how your gonna get it have a really nice outliine, it helps

OpenStudy (anonymous):

I'm going to try doing that from now on. Whenever i see related rates or optimization problems i tend to freak out just because they're a bit lengthy, but the outline would help a lot

OpenStudy (danjs):

That is a good hint to apply the linear approx. "Close to" the number.

OpenStudy (fibonaccichick666):

I teach it for algebra to physics to calc to everything, a game plan helps reduce anxiety

OpenStudy (fibonaccichick666):

and even if you're clueless, you may get a point or two

OpenStudy (danjs):

yes, draw pictures, free body diagrams, givens